OpenStudy (anonymous):

Find the remainder: 1. P(x)=2x^100 divided by (x-1) 2. P(x)=3x^200-5x^100+1 is divided by (x+1) 3. Find the value k so that when P(x)=x^3+3x^2+kx-3 divided by (x-2) the remainder is -17. note: I know how to use synthetic and long division on this but i just can't get the answer. Pls help me how to solve this.

OpenStudy (anonymous):

you can use the divison algorithm to find the remainder, a = bq + r so for number 1 \[2x^{100} = (x-1)q + r\] let x = 1 and solve for r

OpenStudy (anonymous):

division algorithm? okay wait im gonna try this

OpenStudy (anonymous):

ok im confused help pls

OpenStudy (anonymous):

divison algorithm is: Let a be an integer and b be a positive integer. Then there exist unique integers q and r such that a = bq + r, where 0 <= r < b

OpenStudy (anonymous):

so for number 1 you plug in x = 1 you get \[2(1)^{100} = (1-1)q + r\] so you get 2(1) = 0q + r 2 = 0 + r r = 2 so your remainder is 2

OpenStudy (anonymous):

do you understand what i did?

OpenStudy (anonymous):

why is that q was 0?

OpenStudy (anonymous):

0 multiplied by anything is 0

OpenStudy (anonymous):

i let x = 1 since i know 1-1 = 0 and 0q = 0 so i can get r by itself

OpenStudy (anonymous):

so q is equals to 0?

OpenStudy (anonymous):

ohh i get it.. !

OpenStudy (anonymous):

you pretty much want to make q equal to 0 so you can solve for r

OpenStudy (anonymous):

okay one more question in question #3 how can you find the value of k?

OpenStudy (anonymous):

you can use the division algorithm again \[x^3 + 3x^2 + kx - 3 = (x-2)q - 17\] let x = 2 and get: \[2^3 + 3(2^2) + 2k - 3 = (2-2)q - 17 \] so you get 8 + 3(4) + 2k - 3 = 0 -17 8 + 12 + 2k - 3 = -17 20 + 2k - 3 = -17 17 + 2k = -17 2k = 0 k = 0

OpenStudy (ranga):

k = -17

OpenStudy (ranga):

In general, if a polynomial f(x) is divided by (x - a), then the remainder is f(a).

OpenStudy (anonymous):

hm.. im off by a sign

OpenStudy (anonymous):

but k = - 17 is correct

OpenStudy (anonymous):

o wait i see i made a stupid algebraic mistake 17 + 2k = -17 2k = -34 k = -17 sorry haha

OpenStudy (ranga):

happens to all of us :)

OpenStudy (ranga):

For part 2) Find remainder when P(x)=3x^200-5x^100+1 is divided by (x+1) (x+1) = (x -(-1)) so a = -1. P(-1) = 3(-1)^200 - 5(-1)^100 + 1 = ? that will be your remainder.

OpenStudy (anonymous):

Thank u !! I can finally answer this other questions!!

OpenStudy (ranga):

yw.

OpenStudy (anonymous):

your welcome glad to help